# VibeCodeHPC: An Agent-Based Iterative Prompting Auto-Tuner for HPC Code Generation Using LLMs

### v2

@@ IV-B Matrix multiplication
Matrix multiplication is one of the fundamental kernels in scientific computing. This computation is highly compute-intensive; therefore, even when accounting for the data transfer cost between the CPU and GPU, utilizing the GPU – with its high theoretical computational performance – is expected to be advantageous.
@@
Table I shows the performance of the generated code. The technology classification in the tables (OpenMP, MPI, SIMD, CUDA, OpenACC, etc.) corresponds to the directory names where PGs were deployed, representing the parallelization technique each PG handled. Empty cells in the tables indicate that the corresponding parallelization technique was not selected by the agent. In each experiment, the choice of which parallelization technique to use was left to the autonomous judgment of the agents. We can see that the multi-agent configuration achieves the maximum performance in almost all cases. Table II shows the number of generated codes. We see that the multi-agent configuration generates more variations.
@@
Additionally, in EX3, the solo agent violated the requirement definition that fixed the matrix size at 2048, gradually expanding the matrix size to 16384 to improve performance, and ultimately reported achieving 8898 GFLOPS.
@@
Cat Technology Mode Experiments
@@
GPU CUDA Multi 1384.6 3677.2 916.2 1993.0
Solo 804.5 606.9 1097.7 837.0
@@
Cat Technology Mode Experiments
@@ IV-C Himeno Benchmark (Poisson equation solver)
We targeted the “himenoBTM.c” in “Himeno Benchmark 98.” It includes several problem size settings, and we selected SMALL among them. This is too small to offload to the GPU, so CPU utilization is expected. The code includes a problem generation section containing embedded fixed values. To prevent these from being expanded into the computation section and thus defeating the code’s purpose, random perturbations are added to the matrix generation. These perturbations are configured to have no impact on the operation verification. The code displays performance (in GFLOPS) and verifies the correctness of the computation at the end. The correctness is judged when the “gosa” item is 3e-3 or less and the maximum error of all coefficient values is 1e-6 or less.
@@
Table III shows the achieved performance. This case is unsuitable for GPU implementation, and CPU-oriented optimization achieved good performance. Table IV shows the number of generated codes. Figure 6 shows the best performance progression chart created by the SE agent during the multi-agent Himeno benchmark optimization experiment EX3.
Figure 6: Example of best performance progression chart created by SE during multi-agent Himeno benchmark optimization experiment
@@
Cat Tech Mode Experiments
@@
Cat Technology Mode Experiments
@@ IV-D Discussion
Task Execution Stability: Pushing to GitHub and the creation of best performance progression graphs were more stable with the multi-agent configuration. In the solo agent experiments, pushing was not performed in 5 out of 6 experiments, and the best performance progression graphs had insufficient data points.
@@
Quality Assurance through Mutual Monitoring: As mentioned in the matrix multiplication experiment, in cases where the solo agent ignored user instructions and used cuBLAS, the multi-agent configuration automatically detected and prevented this through monitoring by other agents. This mutual monitoring function plays an important role in long-term autonomous operation.
@@ V Conclusion
This study proposed “VibeCodeHPC,” a multi-LLM agent system aimed at automatically tuning HPC programs on supercomputers. This system is designed to specialize a CLI coding agent (Claude Code in this study) for complex HPC programming development by configuring multiple agents with distinct roles that autonomously coordinate their actions. Consequently, VibeCodeHPC not only enables vibe coding with a specialized HPC programming team but also achieves complete automatic tuning through autonomous operation that minimizes user intervention. This paper demonstrates VibeCodeHPC’s most challenging feature – fully autonomous performance optimization – using CPU-oriented matrix multiplication for GPU systems and a Poisson equation solver using Jacobi’s iterative method. This evaluation shows that the multi-agent configuration effectively monitors each other, enabling the generation of more accurate code that closely follows user requirements in a shorter time.
VibeCodeHPC offers a wide range of features, and we hope each one will be evaluated for its effectiveness. However, we wish to emphasize that we have demonstrated its most challenging feature—autonomous operation—which inherently encompasses the evaluation of the effectiveness of its other features. For example, while VibeCodeHPC aims for vibe coding, it allows user intervention at any time through the tmux screen, enabling the incorporation of vibe coding development techniques whenever needed. The advanced autonomous operation achieved by VibeCodeHPC can accelerate development through vibe coding, and with user intervention, VibeCodeHPC can more easily develop the desired program. Furthermore, developing distributed parallel programs is crucial on supercomputers. However, VibeCodeHPC already possesses the capability to handle batch systems and MPI, addressing this within the same fundamental framework as single-node GPU programming. MPI programming capability depends on the underlying Claude Code’s capabilities, but VibeCodeHPC provides the supporting environment for this.
VibeCodeHPC has many directions for expansion. For example, since VibeCodeHPC is an automatic tuning system, integrating existing automatic tuning strategies or tools [20][21][22] with it is expected to enable more efficient automatic tuning. Furthermore, in developing numerical computation code, unlike programs where mere functionality suffices, it is necessary to address challenges specific to numerical computation, particularly numerical error issues. Additionally, mixed-precision approaches are effective for acceleration, necessitating the expansion of features that incorporate such optimization considerations.
The introduction of automatic verification methods such as dependency verification through data flow analysis is also being considered, but there are practical constraints such as execution time limitations. In VibeCodeHPC, strengthening the code-level verification function by audit agents to automatically discover problems that are difficult to detect with user-provided tests alone is a future challenge.

### v3

@@ IV-B Matrix multiplication
Matrix multiplication is a fundamental kernel in scientific computing. This computation is highly compute-intensive; therefore, even accounting for data transfer costs between the CPU and GPU, utilizing GPUs with high theoretical computational performance is expected to be advantageous.
@@
Table I shows the performance of the generated code. The technology classification in the tables (OpenMP, MPI, SIMD, CUDA, OpenACC, etc.) corresponds to the directory names where PGs were deployed, representing the parallelization technique each PG handled. Empty cells in the tables indicate that the corresponding parallelization technique was not selected by the agent. In each experiment, the choice of which parallelization technique to use was left to the autonomous judgment of the agents. We can see that the multi-agent configuration achieves the maximum performance in almost all cases. Table II shows the number of generated codes. We see that the multi-agent configuration generates more variations. Figure 6 shows the best performance progression over time for all experiments in both multi-agent and solo configurations.
@@
Additionally, in EX3, the solo agent violated the requirement definition that fixed the matrix size at 2048, gradually expanding the matrix size to 16384 to improve performance, and ultimately reported achieving 8898 GFLOPS. A similar matrix size violation also occurred in the multi-agent EX2, where the CUDA PG expanded the matrix size to 4096 and 8192 to improve arithmetic intensity, achieving up to 3677 GFLOPS. Although these results were reported in the inter-agent communication logs, neither the SE nor PM agents detected the deviation from the fixed problem size. The performance values in Table I exclude results obtained with non-standard matrix sizes for both cases.
@@
Target Technology Mode Experiments
@@
GPU CUDA Multi 1384.6 1141.0 916.2 1147.3
Solo 804.5 561.7 1097.7 821.3
@@
Target Technology Mode Experiments
@@ IV-C Himeno Benchmark (Poisson equation solver)
We targeted the “himenoBTM.c” in “Himeno Benchmark 98.” It includes several problem size settings, and we selected SMALL ($129\times 65\times 65$ grid, approximately 29 MB of single-precision working data across seven arrays—three multi-component coefficient matrices, pressure, boundary, and two work arrays) with its default iteration count ($\mathrm{NN}=200$). The sub-second execution time enables both configurations to complete many compile–test–evaluate cycles within the time budget. At this scale, both CPU-targeted and GPU-based optimizations are potentially effective, requiring agents to evaluate diverse parallelization strategies—complementing the matrix multiplication experiment where GPU is clearly dominant. This is a conservative setting: the fast turnaround gives the solo agent ample optimization opportunities, whereas larger problem sizes would reduce cycle counts and widen the multi-agent advantage. Evaluation at larger scales remains future work. The original code contains embedded fixed coefficient values (zeros and ones). Without modification, a compiler or an LLM-generated code could apply constant propagation, eliminating the memory-access patterns that the benchmark is designed to measure and producing artificially inflated performance. To prevent this, we add small random perturbations to the coefficient matrices; these perturbations are bounded so as not to affect correctness verification. The code reports performance in GFLOPS and verifies correctness: the residual norm “gosa” must satisfy $\mathrm{gosa}\leq 3\times 10^{-3}$, and the maximum coefficient error must be at most $10^{-6}$.
@@
Table III shows the achieved performance. As expected from the small problem size, CPU-targeted optimizations (OpenMP and MPI) significantly outperformed GPU-based implementations. Notably, multi-agent configurations explored diverse parallelization strategies and achieved the best results via hybrid OpenMP+MPI, whereas solo agents produced fewer non-GPU codes. Table IV shows the number of generated codes. Figure 7 shows the best performance progression over time for all Himeno benchmark experiments in both multi-agent and solo configurations.
Figure 7: Best performance progression for the Himeno benchmark across all experiments. Multi-agent runs (solid lines) consistently outperform solo-agent runs (dashed lines) within the 90-minute time limit.
@@
Target Tech Mode Experiments
@@
Target Technology Mode Experiments
@@ IV-D Discussion
Task Execution Stability: Pushing to GitHub and the creation of best performance progression graphs were more stable with the multi-agent configuration, as shown in Figures 6 and 7. In the solo agent experiments, pushing was not performed in 5 out of 6 experiments, and the best performance progression graphs had insufficient data points.
@@
Quality Assurance through Mutual Monitoring: As mentioned in the matrix multiplication experiment, in cases where the solo agent ignored user instructions and used cuBLAS, the multi-agent configuration automatically detected and prevented this through monitoring by other agents. This mutual monitoring function plays an important role in long-term autonomous operation. However, this monitoring was not comprehensive: in EX2, the CUDA PG’s expansion of the matrix size beyond the specified $n=2048$ went undetected, despite being visible in the ChangeLog and inter-agent communication logs. This suggests that the current monitoring mechanism is more effective at detecting explicit rule violations, such as prohibited library usage, than implicit constraint deviations, such as problem size changes.
@@ V Conclusion
This study proposed VibeCodeHPC, a multi-LLM agent system for automatic tuning of HPC programs on supercomputers. By orchestrating multiple role-specialized agents around a CLI coding agent, VibeCodeHPC enables vibe coding for complex HPC development and achieves fully automated tuning with minimal user intervention. This paper demonstrates this capability using a CPU-targeted matrix multiplication implementation and a Poisson equation solver based on Jacobi’s iterative method for GPU systems, and the evaluation showed that mutual monitoring among agents improves adherence to user requirements and reduces development time.
While VibeCodeHPC provides a wide range of features, this study focuses on demonstrating autonomous operation, the most challenging aspect of the system. Because autonomous operation requires effective integration of other functionalities, its evaluation reflects their combined effectiveness. VibeCodeHPC also allows optional user intervention via the tmux screen, supporting interactive refinement when needed. Furthermore, VibeCodeHPC supports batch job systems and MPI within the same framework as single-node GPU programming, providing the execution and orchestration environment required for distributed parallel development.
VibeCodeHPC has several directions for future expansion. As an automatic tuning system, integrating existing auto-tuning strategies and tools [20][21][22] is expected to further improve tuning efficiency. In addition, extending VibeCodeHPC to address numerical computation–specific challenges, such as numerical error control and mixed-precision optimization, is an important direction. Automatic verification methods, including dependency verification via data flow analysis, are also being considered, although practical constraints such as execution time remain. Strengthening code-level verification through audit agents to detect issues beyond the coverage of user-provided tests is therefore a key future challenge.